OpenStudy (anonymous):

find circumference. arc is 24 cm and angle is 135 degrees

OpenStudy (anonymous):

64

OpenStudy (anonymous):

\[\frac{arc}{circumference}=\frac{135^o}{360^o} \rightarrow circumference =\frac{360 \times arc}{135}\]\[=\frac{360 \times 24}{135}=64cm\]